if (9root 3 + 11 root 2) power 1/3 = root a + root b then prove a = 3 given that a,b are rational

Grade:9

1 Answers

Akshay
8 years ago
9rt3 + 11rt2 = (rt2 + rt3)3,
use a3+b3+3a2b+3ab2=(a+b)3, to prove that